### v3

# Efficient Dataset Distillation via Diffusion-Driven Patch Selection for Improved Generalization

Abstract
Dataset distillation offers an efficient way to reduce memory and computational costs by optimizing a smaller dataset with performance comparable to the full-scale original. However, for large datasets and complex deep networks (e.g., ImageNet-1K with ResNet-101), the extensive optimization space limits performance, reducing its practicality. Recent approaches employ pre-trained diffusion models to generate informative images directly, avoiding pixel-level optimization and achieving notable results. However, these methods often face challenges due to distribution shifts between pre-trained models and target datasets, along with the need for multiple distillation steps across varying settings. To address these issues, we propose a novel framework orthogonal to existing diffusion-based distillation methods, leveraging diffusion models for selection rather than generation. Our method starts by predicting noise generated by the diffusion model based on input images and text prompts (with or without label text), then calculates the corresponding loss for each pair. With the loss differences, we identify distinctive regions of the original images. Additionally, we perform intra-class clustering and ranking on selected patches to maintain diversity constraints. This streamlined framework enables a single-step distillation process, and extensive experiments demonstrate that our approach outperforms state-of-the-art methods across various metrics.
1 Introduction
Figure 1: Comparison of various methods for dataset distillation using diffusion models. (a) Fine-tuning a diffusion model pre-trained on ImageNet-1K to directly generate images. (b) Calculating the clustering center of the original images in latent space to generate images. (c) Identifying the regions in the real images most strongly associated with label text based on the implicit data distribution learned by the diffusion model. The images generated by our method more effectively represent the feature distribution of the original dataset.
The rapid advancement of deep learning has driven impressive performance gains through increasingly deeper and more complex models trained on large-scale datasets [9, 5]. However, training these models demands significant memory and computational resources. Dataset distillation has recently emerged as an effective approach to data compression [17], offering substantial reductions in resource consumption. By optimizing images in pixel space through information matching (e.g., gradient matching), these optimization-based methods produce smaller synthetic datasets that can achieve comparable performance to the original data [37, 36, 1]. Despite these benefits, the vast pixel-level optimization space restricts the feasibility of these methods to small-scale datasets, e.g., CIFAR-10/100 [15], thus limits their generalization.
A recent approach, SRe2[35], scales optimization-based methods to larger datasets (e.g., ImageNet-1K [4]) by using a pre-trained classifier to guide dataset synthesis. However, its strong dependence on proxy models during optimization often hinders the distilled dataset’s generalization, limiting broader applicability. Another critical limitation arises when handling various distillation configurations, such as different subsets of large datasets. Previous optimization-based methods typically incur substantial redundant costs due to repeated distillation requirements. Although recent works [11, 10] aim to address this issue, their applicability remains confined to small-scale datasets.
To address the limitations of current dataset distillation methods, recent approaches [31, 7, 30] that avoid direct image optimization have shown promise for scaling dataset distillation to larger datasets and more complex model architectures. RDED [31] leverages a pre-trained classifier (e.g., ResNet-18 [9]) to select patches randomly cropped from real images and then concatenate them into synthetic images. While RDED achieves substantial performance improvements, it requires careful model selection and retraining of a classifier tailored to the specific dataset. Moreover, concatenated images often lack complete class-relevant features, especially for complex models with higher learning capacity. Minimax [7] addresses distillation by pruning the DiT model [22] to generate synthetic datasets using regularization based on dataset representativeness and diversity. However, Minimax relies on a diffusion model pre-trained on ImageNet-1k [4], requiring fine-tuning for different target datasets, which limits flexibility. D4[30] applies a pre-trained text-to-image Latent Diffusion Model (LDM)[23] to generate synthetic datasets by clustering original images in latent space. However, without constraints on distributional differences between the LDM’s pre-training data (e.g., LAION[26]) and the target dataset (e.g., ImageNet-1K), D4M can produce class-irrelevant or noisy images.
To overcome these limitations, we propose a novel dataset distillation paradigm that leverages diffusion model’s prior to select the most informative image patches. Specifically, for each input image from the original dataset, we first predict the noise map using the pre-trained LDM, with and without the label text prompt, calculating losses for each scenario. The differential loss quantifies representativeness, enabling us to crop patches most relevant to the label guided by the target dataset’s data distribution. Additionally, we can effectively constrain the implicit data distribution learned by the diffusion model by using the pre-trained LDM to identify class-pertinent regions.
To mitigate the inherent diversity of diffusion models, we apply clustering to capture visual features most representative of each class. Specifically, we compute diffusion features for all patches inspired by recent work [32], allowing for efficient clustering and ranking of cropped patches. Fig. 1 illustrates a comparison between our method and previous diffusion-based generative distillation methods. Extensive experiments demonstrate that our approach achieves state-of-the-art (SOTA) performance on large-scale datasets, supporting an efficient, unrestricted one-step distillation process.
Our contributions are summarized as follows:
  • We introduce a novel dataset distillation framework that diverges from conventional methods by leveraging pre-trained diffusion models in an orthogonal way. Our approach uniquely addresses distributional discrepancies between the diffusion model and target dataset, offering a new perspective on improving dataset fidelity and applicability.
  • Our proposed paradigm enables a highly efficient, one-step distillation process, eliminating the need for model training and fine-tuning. Additionally, it avoids repetitive distillation across subclass combinations or compression ratios, which substantially broadens the potential applications and flexibility of dataset distillation.
  • Extensive experiments demonstrate that our method consistently surpasses existing training-free approaches, particularly on large-scale datasets, underscoring the effectiveness and robustness of our framework in various challenging settings.
2 Related Work
Figure 2: (a): Stage I leverages image and corresponding differential class text prompts as inputs to the diffusion model to select the regions that most represent class-relevant features. (b): Stage II first computes DIFT of the patches and performs clustering to aggregate the most representative visual elements for each class. The evaluation process utilizes soft labels provided by a pre-trained teacher model. Rather than existing methods employing diffusion models, our proposed method maximizes the restoration of the original data distribution.
2.1 Dataset Distillation
Dataset distillation was initially regarded as a meta-learning problem [34]. It involves minimizing the loss function of the synthetic dataset through a model trained on this dataset. Since then, several approaches have been proposed to enhance the performance of dataset distillation. One category of methods utilizes ridge regression models to approximate the inner loop optimization [20, 21, 41]. To address the issue of data distribution bias in meta-learning, another category of methods selects an information space and computes proxy information to align the synthetic dataset with the original dataset. DC [39] and DSA [37] match the weight gradients of models trained on the real and synthetic dataset, respectively, while DM [38], CAFE [33] and DataDAM [25] focus on using feature distance between the two datasets as matching targets. MTT [1] and TESLA [3] match the model parameter trajectories obtained from training on the real dataset and the synthetic dataset.
Recent methods propose the decoupled distillation strategy and successfully extend the dataset distillation application to large-scale datasets. SRe2[35] leverages a pre-trained classifier to compute the cross-entropy loss of the synthetic dataset for updates, while RDED [31] calculates the confidence of randomly cropped patches from the original images with a pre-trained classifier and concatenates them together to form the synthetic dataset. Recent methods [2, 40] have introduced GAN as a parameterization technique. While with the development of diffusion models, Minimax [7] and D4[30] introduce the diffusion models to directly generate synthetic datasets, achieving promising performance. However, existing diffusion-based methods are constrained by the implicit data distribution learned by diffusion models, leading to significant distributional bias in the synthetic dataset. In addition, aforementioned methods still fail to complete the distillation process in one pass for all settings. To address these issues, we utilize diffusion model to actively identify the most class-relevant regions in the original images, enabling an efficient dataset distillation process.
2.2 Diffusion Model
As a class of generative approaches, diffusion-based frameworks learn to progressively convert Gaussian-distributed noise vectors $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into samples matching target distributions $\mathcal{X}$, demonstrating remarkable effectiveness across diverse applications [42, 43, 28]. These models employ a neural network $\epsilon_{\theta}(\mathbf{x},t)$ parameterized by $\theta$ to estimate noise components, operating through sequential refinement steps denoted by a temporal parameter $t$. The denoising procedure typically involves iterative updates where the network processes corrupted inputs $\mathbf{x}$ at progressive noise levels determined by $t$.
3 Method
In this section, we delve into the specifics of our method, aiming to address the limitations of previous methods. The overall framework of our method is shown in Fig. 2.
3.1 Preliminaries
The training protocol for latent diffusion models involves two core phases. Initially, an encoder network $E(\cdot)$ compresses input samples $\mathbf{x}$ into latent representations $\mathbf{z}=E(\mathbf{x})$. These latent codes undergo systematic corruption through a diffusion trajectory determined by uniformly sampled timesteps $t$, implemented via the forward process formulation:
$\text{noise}(\mathbf{z},\epsilon,t)=\sqrt{\bar{a}{t}}\mathbf{z}+(1-\sqrt{\bar{a}{t}})\epsilon,$ (1)
where time-dependent attenuation factors $\sqrt{\bar{a}_{t}}$ govern the noise blending proportions throughout the diffusion stages.
The denoising network $\epsilon_{\theta}$ learns to recover original signals by processing corrupted latent representations paired with their corresponding timestep indices. This is achieved through the optimization objective:
$\mathcal{L}_{t}(\mathbf{z},\epsilon)=|\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),t)-\epsilon|_{2}^{2}.$ (2)
During the generation phase, novel samples are synthesized through a reverse diffusion process [29, 13] that progressively refines Gaussian noise samples $\mathbf{z}_{T}$ over multiple iterations. For conditional generation tasks like text-to-image synthesis, the framework incorporates semantic conditioning signals $c$ (e.g., text embeddings) as auxiliary inputs. This extends the denoiser architecture to $\epsilon_{\theta}(\mathbf{z},t,c)$, modifying the training objective accordingly:
$\mathcal{L}_{t}(\mathbf{z},\epsilon,c)=|\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),t,c)-\epsilon|_{2}^{2}.$ (3)
Algorithm 1 Pseudocode of the proposed method
0:  ($\mathcal{T}$, $\mathcal{Y}$): Real dataset and corresponding label texts.
0:  $\mathcal{P}$: Selected patches.
0:  $I$: Pre-trained image encoder.
0:  $T$: Pre-trained text encoder.
0:  $\mathcal{U}$: Pre-trained time-conditional U-Net.
0:  $N$: Number of diffusion time-steps.
0:  $C$: Number of top cluster centers.
1:  initialize $\mathcal{S}=\emptyset$
2:  $Z=I(\mathcal{T})\sim P_{z}$
3:  for  each class $Y\in\mathcal{Y}$ do
4:     $c=T(Y)$
5:     initialize $\mathcal{S}_{Y},\mathcal{P}_{Y}=\emptyset$
6:     for  each latent $\mathbf{z}\in Z$ do
[…]
8:           $t\sim U(0.1,0.7)$
9:           $L_{i}=L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)$
10:        end for
11:        $R(\mathbf{z}|c)=\frac{1}{N}\sum_{i}L_{i}$
12:        $\mathcal{P}_{Y}\leftarrow\mathcal{P}_{Y}\cup\mathop{\arg\max}\limits_{\mathbf{p}\in\mathbf{z}}R(\mathbf{p}|c)$
13:     end for
14:     $cluster(\mathcal{U}(I(\mathcal{P}_{Y}),c,1.6))$
[…]
16:        $\mathcal{S}_{Y}\leftarrow\mathcal{S}_{Y}\cup top(\mathcal{P}_{Y}^{i})$
17:     end for
18:     $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{Y}$
19:  end for
19:  $\mathcal{S}$: Distilled dataset.
3.2 Optimal Class-relevant Feature Selection
In recent years, several methods propose the utilization of pre-trained diffusion models to directly generate images in order to improve dataset distillation performance on large-scale datasets. However, these methods often face the challenge of distributional differences between the pre-training dataset and the target dataset. Minimax [7] relies on fine-tuning a diffusion model pre-trained on ImageNet-1K, while D4[30] attempts to obtain representative samples through clustering in latent space. Both of them remain constrained by the distributional shifts introduced during the unconstrained denoising process.
To address the issue of data heterogeneity, we introduce a framework consisting of two stages that utilizing the diffusion model for localization rather than generation. As illustrated in the Stage I of Fig. 2, we first utilizes diffusion models as a zero-shot image-wise classifiers similar to recent work [18], which can be formulated as:
$p(c_{i}|\mathbf{z})$ $=\frac{\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),\;t,\;c_{i})-\epsilon]\}}{\sum_{j}\mathrm{exp}\{-\mathbb{E}_{\epsilon,t}[\epsilon_{\theta}(\text{noise}(\mathbf{z},\epsilon,t),\;t,\;c_{j})-\epsilon]\}}$ (4)
$=\frac{1}{\sum_{j}\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c_{i})-L_{t}(\mathbf{z},\epsilon,c_{j})]\}}.$
where $c_{i}\in[C]$ and $[C]$ represents the label space, $\mathbf{z}$ is the latent of input image $\mathbf{x}$. However, such an approach can only select the image with the highest classification probability, lacking further constraints. Additionally, a larger class space can lead to significant computational overhead. Using the classifier-free guidance[12], we redefine the label space for a specific label $c$ and rewrite Eq. 4 as follow:
$p(c|\mathbf{z})=\frac{1}{1+\mathrm{exp}\{\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)]\}},$ (5)
where $\phi$ and $c$ represents the classifier-free guidance and label text prompt respectively (i.e., ”An image of .” and ”An image of c”). To simplify computation, we define the representative of an image as follow:
$R(\mathbf{z}|c)=\mathbb{E}_{\epsilon,t}[L_{t}(\mathbf{z},\epsilon,c)-L_{t}(\mathbf{z},\epsilon,\phi)].$ (6)
Obviously Eq. 5 and Eq. 6 are monotonically equivalent. Based on this, to further reduce the complexity of the dataset, we calculate the $R(\mathbf{p}|c)$ within each patch $\mathbf{p}$ in image and select the most typical patches as the synthetic dataset. The detailed workflow is depicted in Algorithm 1.
3.3 Visual Elements Aggregation
Although we have successfully selected the most representative patches, the overly complex data often introduce excessive randomness. To enhance the representative and reduce the randomness of the synthetic dataset, we perform an effective selecting strategy on the cropped patches as shown in the Stage II of Fig. 2. With diffusion model, we select DIFT [32] as the feature embedding for applying clustering, where noise at a specific time step is added to the latent of input image, and passed through a U-Net [24] with text condition to obtain the intermediate layer activations as features. On this basis, we use k-means [19] to cluster the selected patches. To better leverage clustering information, we have designed a ranking rule: 1) Intra-cluster. patches are ranked based on their distance to the centroid instead of corresponding $R(\mathbf{p}|c)$. 2) Inter-cluster. clusters are ranked based on the median $R(\mathbf{p}|c)$ of the patches within each cluster, the partial visualization is shown in Fig. 3. In practice, we fix the number of cluster centers at 32 and prioritize the selection of patches from the top 10 clusters under all compression settings to ensure both representativeness and diversity. A noteworthy point is that, unlike D4M, which dynamically updates cluster centers and generates samples accordingly, we utilize static clustering to obtain more class-relevant visual representations.
3.4 Fine-grained Image Reconstruction
When constructing the synthetic dataset for large-scale datasets (e.g., ImageNet-1K) using patches, we observed the following phenomenon: when the synthetic dataset size is limited with smaller IPC, the model’s ability to learn complex representations is constrained, thus the samples in synthetic dataset must represent the most prevalent patterns. By focusing on learning from these representative features, the model can capture essential discriminative information across classes. Therefore, when IPC $\leq 10$, we follow the experimental setup of RDED by concatenating multiple patches into a single image, allowing the model to focus on representative class features as shown in Fig. 4. However, as the IPC increases, complete feature is more likely to provide representations that closely align with statistical patterns, while excessively fragmented features tend to hinder network’s ability to improve accuracy and generalization, especially in deeper models with stronger learning capacity. Thus, we adopt the experimental settings of Minimax and D4M, where each patch is treated as an individual image. Experimental results demonstrate that our strategy yields effective outcomes.
Figure 3: Visualization of clustering. The top five patches from the top three clusters within the golden retriever class. Each row represents a single cluster.
3.5 Effective Label Calibration
Based on previous research [35, 31, 30] and preliminary experiments, we find that training large-scale synthetic datasets with hard labels can hinder the model from accurately learning useful information. In contrast, leveraging soft labels provided by the teacher model for the corresponding synthetic datasets can significantly improve the accuracy and generalization of the student model. Therefore, for a fair comparison, we also adopt FKD [27] which is utilized in RDED and D4M to align soft label during the training process, which can be formulated as:
$\mathbf{\theta}^{t+1}=\mathop{\arg\min}\limits_{\mathbf{\theta}\in\Theta}L_{KL}(T^{t}(\mathbf{x}),S_{\theta}^{t}(\mathbf{x})),$ (7)
where $T^{t}(x)$ and $S^{t}(x)$ represents the prediction for the distilled image $\mathbf{x}$ at training epoch $t$, $L_{KL}$ represents the Kullback–Leibler (KL) divergence.
Figure 4: Visualization comparison between RDED and ours under IPC=10. Our method can provide more complete class-specific representation.
Figure 5: Visualization comparison of images from the top ten classes of the ImageNet-1k dataset is presented (not cherry picked). From top to bottom, each row corresponds to the real dataset, RDED, Minimax, D4M, and ours under IPC=50. Compared to RDED, our method provides more comprehensive class-specific features. In contrast to the approach of directly using diffusion to generate synthetic datasets, our method fully adheres to the distribution of the real dataset and demonstrates a sharper focus on class-specific features.
ResNet-18 ResNet-101
Dataset IPC SRe2L Minimax D4M RDED OursC OursS SRe2L Minimax D4M RDED OursC OursS
10 29.4 $\pm$ 3.0 57.2 $\pm$ 0.8 57.4 $\pm$ 0.4 61.4 $\pm$ 0.4 62.5 $\pm$ 0.2 60.4 $\pm$ 0.3 23.4 $\pm$ 0.8 49.2 $\pm$ 0.3 49.6 $\pm$ 0.6 54.0 $\pm$ 0.4 56.4 $\pm$ 0.3 52.1 $\pm$ 0.4
ImageNette 50 40.9 $\pm$ 0.3 84.4 $\pm$ 0.5 84.8 $\pm$ 0.2 80.4 $\pm$ 0.4 83.4 $\pm$ 0.5 85.8 $\pm$ 0.4 36.5 $\pm$ 0.7 80.4 $\pm$ 0.9 80.6$\pm$ 0.4 75.0 $\pm$ 1.2 78.1 $\pm$ 0.4 82.4 $\pm$ 0.5
100 50.2 $\pm$ 0.4 90.1 $\pm$ 0.6 90.4 $\pm$ 0.7 89.6 $\pm$ 1.0 90.2 $\pm$ 0.2 91.6 $\pm$ 0.4 47.8 $\pm$ 0.6 88.7 $\pm$ 0.6 88.2 $\pm$ 1.4 88.2 $\pm$ 1.1 88.8 $\pm$ 0.2 90.4 $\pm$ 0.6
10 20.2 $\pm$ 0.2 38.1 $\pm$ 0.6 36.8 $\pm$ 1.2 38.5 $\pm$ 2.1 44.3 $\pm$ 0.5 40.1 $\pm$ 0.5 17.7 $\pm$ 0.9 34.6 $\pm$ 0.5 31.8 $\pm$ 1.2 31.3 $\pm$ 1.3 37.6 $\pm$ 0.4 34.8 $\pm$ 0.2
ImageWoof 50 23.3 $\pm$ 0.3 71.2 $\pm$ 0.6 71.4 $\pm$ 1.4 68.5 $\pm$ 0.7 70.1 $\pm$ 0.5 73.5 $\pm$ 1.4 21.2 $\pm$ 0.2 66.4 $\pm$ 0.6 67.2 $\pm$ 0.8 59.1 $\pm$ 0.7 65.2 $\pm$ 0.4 70.6 $\pm$ 0.2
100 37.6 $\pm$ 0.3 77.4 $\pm$ 0.2 78.5 $\pm$ 0.3 77.2 $\pm$ 0.6 76.9 $\pm$ 0.6 80.2 $\pm$ 0.8 34.4 $\pm$ 0.7 74.0 $\pm$ 0.4 73.4 $\pm$ 0.8 74.1 $\pm$ 0.5 73.8 $\pm$ 0.2 76.8 $\pm$ 0.4
10 9.5 $\pm$ 0.4 31.6 $\pm$ 0.7 33.4 $\pm$ 0.2 36.0 $\pm$ 0.3 37.0 $\pm$ 0.3 35.1 $\pm$ 0.3 6.4 $\pm$ 0.1 30.1 $\pm$ 0.7 33.1 $\pm$ 0.7 33.9 $\pm$ 0.1 36.5 $\pm$ 0.6 33.7 $\pm$ 0.4
ImageNet-100 50 27.0 $\pm$ 0.4 64.0 $\pm$ 0.5 62.7 $\pm$ 0.6 61.6 $\pm$ 0.1 64.2 $\pm$ 0.3 68.8 $\pm$ 0.6 25.7 $\pm$ 0.3 64.8 $\pm$ 0.6 65.3 $\pm$ 0.6 66.0 $\pm$ 0.6 67.3 $\pm$ 0.1 69.4 $\pm$ 0.6
100 39.4 $\pm$ 0.1 76.4 $\pm$ 0.3 76.8 $\pm$ 0.5 75.2 $\pm$ 0.3 75.2 $\pm$ 0.4 77.5 $\pm$ 0.2 38.2 $\pm$ 0.5 78.1 $\pm$ 0.7 77.6 $\pm$ 0.7 77.2 $\pm$ 0.3 76.9 $\pm$ 0.5 79.4 $\pm$ 0.6
10 21.3 $\pm$ 0.6 44.3 $\pm$ 0.3 41.9 $\pm$ 0.3 42.0 $\pm$ 0.1 43.9 $\pm$ 0.2 42.1 $\pm$ 0.3 30.9 $\pm$ 0.1 46.6 $\pm$ 0.7 44.6 $\pm$ 0.7 48.3 $\pm$ 1.0 51.2 $\pm$ 0.4 47.4 $\pm$ 0.3
ImageNet-1K 50 46.8 $\pm$ 0.2 56.6 $\pm$ 0.9 55.9 $\pm$ 0.3 56.5 $\pm$ 0.1 58.1 $\pm$ 0.3 59.4 $\pm$ 0.2 60.8 $\pm$ 0.5 62.6 $\pm$ 0.6 63.4 $\pm$ 1.0 61.2 $\pm$ 0.4 62.7 $\pm$ 0.2 65.4 $\pm$ 0.7
100 52.5 $\pm$ 0.5 58.6 $\pm$ 0.3 59.6 $\pm$ 0.4 60.5 $\pm$ 0.8 60.7 $\pm$ 0.3 61.8 $\pm$ 0.5 65.4 $\pm$ 0.2 67.3 $\pm$ 0.5 66.1 $\pm$ 0.9 65.4 $\pm$ 1.3 67.7 $\pm$ 0.1 70.0 $\pm$ 0.3
Table 1: Performance comparison with methods applicable to large-scale datasets. Following the evaluation protocol of RDED, we use ResNet-18 as the teacher model for generating the dataset and evaluate on ResNet-18/101. OursC and OursS represent that each image in the synthetic datasets is concatenated patches and a single patch, respectively.
ConvNet ResNet-18
Dataset IPC DSA IDM TESLA DATM D4M Ours SRe2L RDED Ours
CIFAR10 10 52.1 $\pm$ 0.5 58.6 $\pm$ 0.1 66.4 $\pm$ 0.8 66.8 $\pm$ 0.2 56.2 $\pm$ 0.4 48.8 $\pm$ 0.4 29.3 $\pm$ 0.5 37.1 $\pm$ 0.3 36.4 $\pm$ 0.2
50 60.6 $\pm$ 0.5 67.5 $\pm$ 0.1 72.6 $\pm$ 0.7 76.1 $\pm$ 0.3 72.8 $\pm$ 0.5 66.2 $\pm$ 0.3 45.0 $\pm$ 0.7 62.1 $\pm$ 0.1 61.0 $\pm$ 0.6
CIFAR100 10 32.3 $\pm$ 0.3 45.1 $\pm$ 0.1 41.7 $\pm$ 0.3 47.2 $\pm$ 0.4 45.0 $\pm$ 0.1 47.3 $\pm$ 0.4 27.0 $\pm$ 0.4 42.6 $\pm$ 0.2 41.5 $\pm$ 0.4
50 42.8 $\pm$ 0.4 50.0 $\pm$ 0.2 47.9 $\pm$ 0.3 55.0 $\pm$ 0.2 48.8 $\pm$ 0.3 57.6 $\pm$ 0.2 50.2 $\pm$ 0.4 62.6 $\pm$ 0.1 63.8 $\pm$ 0.3
TinyImageNet 10 - 21.9 $\pm$ 0.3 - 31.1 $\pm$ 0.3 35.4 $\pm$ 0.2 40.3 $\pm$ 0.2 16.1 $\pm$ 0.2 41.9 $\pm$ 0.2 40.2 $\pm$ 0.5
50 - 27.7 $\pm$ 0.3 - 39.7 $\pm$ 0.3 41.2 $\pm$ 0.5 49.2 $\pm$ 0.7 41.1 $\pm$ 0.4 58.2 $\pm$ 0.1 58.5 $\pm$ 0.2
Table 2: Performance comparison on low-resolution datasets. To compare with information matching distillation methods, we use ConvNet-3 and ConvNet-4 as the backbone for CIFAR10/100 and TinyImagenet respectively. For decoupled distillation methods, we use ResNet-18 as the backbone. All the performance are evaluated on the same model architecture.
4 Experiments
4.1 Experimental Settings
Datasets and Architectures.
To evaluate our proposed method on low-resolution datasets, we conduct experiments on CIFAR-10/100 [15] and TinyImageNet [16], with ConvNet [6] serving as the backbone. For high-resolution data, we perform evaluations on ImageNet-1K [4] and its widely used subsets, including ImageNet-100 [14], ImageNette and ImageWoof [1], utilizing ResNet-18 [9] as the backbone.
Baselines.
In addition to data-matching methods that are effective on low-resolution data: DSA [37], CAFE [33], and DATM [8], we consider recent methods applicable to large-scale datasets as strong competitors.
  • SRe2[35], by updating the synthetic dataset using the CE-loss of a pre-trained classifier and introducing soft label matching, was the first method to be feasibly applied to large-scale datasets.
  • RDED [31] significantly improved computational efficiency and performance by evaluating randomly cropped patches with a pre-trained classifier and concatenating them together to generate images.
  • Minimax [7] proposed fine-tuning DiT model pre-trained on ImageNet-1K by using the distribution of target dataset as a regularization to directly generate images.
  • D4[30] extended the range of diffusion models to pre-trained text-to-image Stable Diffusion, utilizing the cluster centers of latent to generate synthetic datasets.
Evaluation.
In line with previous work, we set IPC=10 and 50 for low-resolution datasets, and IPC=10, 50, and 100 for high-resolution datasets to enable more efficient learning. During the training process, we employ a fixed pre-trained classifier to provide soft labels. Subsequently, five randomly initialized networks are trained on the distilled dataset, with the average performance of these networks evaluated.
Implementation details.
When calculating $R(x|c)$ of images, we select a time-step range of [0.1, 0.7], whereas for the computation of DIFT, a time-step of 1.60 is employed. Across all IPC settings, we prioritize cropping the patches with resolution of 224x224 and selecting them from the top 10 ranked clusters after clustering. All the experiments were carried out using PyTorch on NVIDIA RTX-3090 GPUs.
4.2 Performance Improvements
High-resolution results.
For high-resolution datasets, we first resize the smallest edge of images to 256 pixels while maintaining height-width ratio. The resized images are then fed into the diffusion model to calculate $R(x|c)$ of the images and that of all patches with a 224x224 averaging pool kernel. As aforementioned, we adopt two settings for different compressing ratio to enhance the image construction performance. Specifically, for IPC $\leq$ 10, we select the top 4 patches from the top 10 clusters, resize each to 112x112, and concatenate them into a single image. When IPC $\geq$ 20, we directly use patches without resizing operation. For instance, when IPC = 100, we select 10 patches from the top ten clusters. If a cluster contains insufficient patches, we supplement the synthetic dataset by selecting the highest-scoring patches. As shown in Tab. 1, our proposed method demonstrates significant improvements across various datasets and IPC settings. In specific configurations, our approach exceeds baseline’s performance by more than 10%, further validating the superiority of our method.
Low-resolution results.
For low-resolution datasets, we maintain the same setting as in RDED, utilizing the diffusion model to calculate the $R(x|c)$ of full-image without cropping, followed by clustering. We utilize ConvNet-3 and ConvNet-4 to generate soft labels for CIFAR-10/100 and TinyImageNet, respectively, and evaluate the synthetic datasets on the same architecture. As demonstrated in Tab. 2, the incorporation of the diffusion model yields performance on CIFAR-10 that is comparable to, or even exceeds, that of existing methods when applied within our proposed framework. On more complex datasets, such as CIFAR-100 and TinyImageNet, our approach outperforms all current optimization-based methods. Under high IPC settings, our method achieves SOTA performance, surpassing all prior techniques.
Efficient one-step distillation.
Distinct from all previous methods, our proposed method achieves a fully unrestricted distillation process and demonstrates SOTA performance. The limitations of existing dataset distillation methods are outlined as follows, and Tab. 3 represents a comprehensive comparison:
  • Large-scale datasets. Optimization-based dataset distillation methods using information matching often suffer from excessive computational overhead, rendering them impractical for large-scale datasets and leading to sub-optimal performance.
  • Zero-shot. Relying on pre-trained classifiers require selecting an appropriate model architecture and training it for the target dataset, the nasty teachers even disrupt the distillation process. Although Minimax attempts to address this issue through the use of diffusion model, it is still constrained by DiT trained on ImageNet-1K.
  • Class-combination. Repeated distillation processes are required when dealing with different class-combination. For example, in ImageNet-1K, a 10-class classification task can result in $C_{1000}^{10}$ possible combinations, leading to significant computational waste.
  • IPC. Similarly, varying IPC settings necessitate multiple distillation, also causing substantial time overhead. While D4M addresses the class-combination issue, it continues to necessitate recalculating different numbers of cluster centers with different IPC settings.
Method Large-scale Zero-shot Class-combination IPC
DATM - - - -
SRe2L - -
Minimax - -
D4M -
RDED -
Ours
Table 3: We present an analysis of the generalization capabilities of various distillation methods under different settings, where ✓denote “Satisfactory”.
Teacher Network Student Network
ResNet-18 ResNet-50 ResNet-101 MobileNet-V2 EfficientNet-B0 Swin-T ViT-B
ResNet-18 59.4 62.7 65.4 53.4 60.1 62.1 69.2
ResNet-50 52.2 64.8 66.7 48.9 56.2 51.3 60.1
ResNet-101 50.1 63.7 66.5 46.1 55.5 42.8 50.5
MobileNet-V2 55.2 62.5 65.0 55.3 60.5 57.2 62.3
EfficientNet-B0 45.8 59.1 61.1 45.0 58.4 51.7 60.3
Swin-T 34.4 51.2 52.7 35.1 44.1 51.1 50.2
ViT-B 31.2 46.9 49.3 32.7 40.0 41.3 46.2
Table 4: Synthetic dataset performance on ImageNet-1K with various teacher-student architectures under IPC=50. Our proposed method shows significant improvement with both CNNs and ViTs. “ ” denotes the best performance of student network.
4.3 Cross-architecture Generalization
Following previous work, we evaluate the performance of the synthetic dataset using different teacher-student combinations. It can be observed that, when using similar network architectures as teacher-student pairs (e.g., both CNNs), shallower networks generally yield better performance as teacher models, while the larger networks show stronger learning ability and obtain the best performance when the size of synthetic datasets increases. We further compare the performance of the synthetic dataset on extremely different network architectures, e.g., CNNs and ViTs. As a student network, the ViT-based networks leverages its global attention mechanism to attain the comparable performance with more realistic images. However, as a teacher network, ViT-based networks does not have strong ability to teach the student networks, yielding weak performance.
4.4 Ablation Study
Ablation IPC-10 IPC-50 IPC-100
Dataset: ImageWoof
w/o Cluster 41.8 70.4 79.1
w/ Cluster-CLIP 39.8 69.5 79.5
w/ Cluster-DIFT 44.3(+2.5) 73.5(+3.1) 80.2(+1.1)
Dataset: ImageNet-1K
w/o Cluster 42.4 58.1 61.0
w/ Cluster-CLIP 42.1 59.0 60.7
w/ Cluster-DIFT 43.9(+1.5) 59.4(+1.3) 61.8(+0.8)
Table 5: Performance Comparison on different data selecting strategys. Cluster-CLIP and Cluster-DIFT represent clustering with CLIP feature and DIFT respectively.
Clustering strategy.
To investigate the effectiveness of the two main stages of our proposed method, we conduct an ablation study to investigate whether clustering operations enhance the representative of synthetic datasets. Moreover, we conduct experiments to explore the differences between using CLIP features and DIFT during the clustering process. The experimental results as shown in Tab. 5, demonstrates that the performance of the synthetic dataset improves further after incorporating the clustering operation. This indicates that reducing sample randomness enables the model to learn more representative information. Additionally, we found DIFT more effective for achieving clustering operations compared to CLIP feature.
Real Images number.
Although performing the distillation operation once on the entire dataset offers the potential for one-step processing, the substantial time overhead makes this approach difficult to apply in practice. To address this issue, we explore the effect of the number of original images used for key patch extraction on performance. As shown in Fig. 6, we randomly select images as a subset of the original dataset and evaluate the corresponding performance. Given that random selection is an unbiased data sampling method, our proposed achieves performance comparable to others, even when the number of original samples is limited. When 300 images are randomly selected from original dataset as input, the performance is nearly on par with the full dataset, resulting in a 4x speedup for the ImageNet-1K dataset.
Figure 6: Ablation on number of real images used.
5 Conclusion
We propose a novel approach that operates orthogonally to existing diffusion-based generative dataset distillation methods, addressing the challenge of distribution mismatch between pre-trained diffusion models and target datasets. Through a two-stage framework, our method achieves a highly efficient, unrestricted one-step distillation process, eliminating the need for repeated distillations under varied settings. Extensive experiments demonstrate that our method significantly improves performance on large-scale datasets and deeper models, while achieving competitive results on low-resolution datasets. Additionally, our work provides a fresh perspective on leveraging diffusion models for dataset distillation.
Supplementary Material
1 Training Efficiency
For a more comprehensive comparison, we evaluate the synthesis time per image and GPU memory consumption of methods that can scale on large-scale datasets. Since previous works on different settings (e.g., different class number of target dataset) require repeated distillation, we measure the time required to generate a single image on the ImageNet-1K under IPC=100 to ensure the fairness. The experimental results are shown in Tab. 6, our method not only achieves the best results but also maintains training efficiency.
SRe2L Minimax D4M RDED Ours
Time (h) 106 256 52 1 95
Accuracy (%) 52.8 58.8 59.3 60.5 61.8
Table 6: Synthesis time per image, GPU memory consumption and corresponding accuracy on ImageNet-1K under IPC=50.
2 Diffusion time intervals
To calculate the difference in predicted loss corresponding to UNet when using different prompts (i.e., w. and w.o. label text), multiple diffusion time steps $t$ are sampled as time conditions and averaged. Therefore, we investigated whether varying the range of time step values would impact the performance of the synthetic dataset.
As shown in Tab. 7, we selected different minimum and maximum edge as the range for the time steps and evaluated the corresponding performance of the synthetic dataset. It can be observed that the dataset achieves the best performance when using intermediate time steps and the small time steps are essential, which is consistent with the conclusions drawn in recent works [18] [32]. In practical applications, considering the generality of ImageNet-1K, we applied the same range of diffusion time steps to all of its subsets (e.g., ImageWoof).
End
Start 0.1 0.3 0.5 0.7 0.9
0.1 - 58.5 58.8 59.4 59.2
0.3 - - 58.1 58.4 59.0
0.5 - - 57.5 58.1 58.6
0.7 - - - - 57.2
Table 7: Ablation study of the range of time steps on ImageNet-1K under IPC=50.
3 Time steps on the computation of DIFT
An important component of our proposed framework involves calculating the DIFT [32] for all patches within the same class and using it as the feature for clustering. The process of calculating DIFT first requires performing a diffusion operation on the input patch at specific time steps. We conduct ablation experiments to investigate the impact of using different time steps to calculate DIFT on the performance of the synthetic dataset.
The experimental results are shown in Fig. 7. As observed, using earlier time steps yields better performance. We hypothesis this is because the earlier time steps of the diffusion model focus more on the low-frequency features of the image, which often provide more effective information for classification tasks. In practical applications, we also follow the results from ImageNet-1k and choose $t=1.60$ as the general setting.
Figure 7: Ablation study of time steps used to calculate the predicted noise under IPC=50.
Teacher Network Student Network
ResNet-18 ResNet-50 ResNet-101 MobileNet-V2 EfficientNet-B0 Swin-T ViT-B
ResNet-18 43.9 50.1 51.2 36.1 46.1 31.2 22.3
ResNet-50 34.6 44.0 42.2 27.8 38.1 27.4 19.8
ResNet-101 33.2 41.4 42.5 28.2 37.0 26.1 17.0
MobileNet-V2 40.1 48.3 46.7 37.5 46.0 30.1 20.5
EfficientNet-B0 29.8 38.4 38.7 27.4 38.9 30.3 21.3
Swin-T 18.9 30.2 31.4 21.6 31.7 26.7 16.0
ViT-B 16.8 27.8 28.9 19.2 28.9 23.6 15.0
Table 8: Synthetic dataset performance on ImageNet-1K with various teacher-student architectures under IPC=10.
Figure 8: Ablation study of the number of cluster center M and the number of selected cluster N on ImageNet-1K.
4 Generalization Ability
To better investigate the generalization capability of our proposed method under different settings, we evaluated its performance using various teacher-student model pairs at IPC=10. Unlike the IPC=50 setting, each image generated is composed of four patches.
As shown in Tab. 8, our proposed method still demonstrates strong performance on CNNs. Additionally, it can be observed that ResNet18 is the most suitable architecture to be used as a teacher model when the sample size is smaller. However, for ViTs, due to their low-pass filter characteristics, effective learning becomes challenging when the sample size is limited and class-specific features are incomplete. The best performance achieved when using DiT as a student model is only 50% of that of the CNN student model under the same settings.
5 Number of cluster centers
To enhance the representativeness of the synthetic dataset, we performed clustering on all patches to extract the most class-relevant visual elements. To gain deeper insights into the impact of the clustering operation on the performance of the synthetic dataset, we conduct additional ablation experiments focusing on the number of cluster centers. Specifically, we evaluate the performance of the synthetic dataset using different numbers of cluster centers $M$, as well as the number $N$ of top clusters used for extracting patches after clustering. In practical applications, when IPC $\leq$ 10, the number of patches provided by each cluster is 4*IPC // $N$, otherwise the number is IPC // N. The remaining patches are selected sequentially from the unused ones based on $R(p|c)$ of each patch $p$ in descending order.
The experimental results are shown in Fig. 8, indicating that when $N$ is fixed, both too few or too many clusters lead to a failure in balancing diversity and representativeness in the synthetic dataset, resulting in a decline in performance. A similar trend was observed in the ablation experiment where $M$ was fixed and $N$ varied.
Patch Size IPC-10 IPC-50 IPC-100
Image Size: 128x128
112x112 42.6 70.2 77.8
Image Size: 256x256
112x112 40.4 67.4 76.8
224x224 44.3 73.5 80.2
Image Size: 512x512
112x112 38.6 65.3 71.4
224x224 43.1 69.5 78.4
448x448 45.1 73.2 81.0
Table 9: Ablation study of different size of real images and patches on ImageWoof
6 Impact of Resizing
When using diffusion to process images, the image size often has a certain degree of impact. On the other hand, when calculating $R(p|c)$ for each patch, we used a fixed-size average pooling kernel to obtain the mean value of the patch. Therefore, we designed an ablation experiment to explore whether different input image sizes and patch sizes significantly affect the performance of the synthetic dataset. In practical applications, we proportionally scale the original image such that its size of shortest edge matches the selected size. For different patch sizes, under the setting of IPC=10, we scale all patches to 112x112 in order to concatenate 4 patches into a 224x224 image. Under the setting of IPC=50, we scale all patches to 224x224.
The experimental results, as shown in Tab. 9, indicate that when the original image size closely matches the standard output size of SD v1.5, better performance are achieved. Additionally, when the size of the cropped patches is closer to the scaled original image size, the image tends to exhibit more complete class-relevant features, leading to better performance. Consider the balance between performance and efficiency, We choose 256x256 as the size of the real images, and 224x224 as the size of the patches. Comparison of visualization with different crop sizes when using 256x256 as the size of the real images are shown in Fig. 9, An excessive size difference between patches and the original image can cause the generated dataset to overly focus on low-frequency features (e.g., noses of dogs), leading to performance degradation.
Figure 9: Visualizations of the patches with different cropped size within class “Golden Retriever”.
7 Qualitative Analysis
To better understand the superiority of the datasets generated by our proposed method, we used t-SNE to visualize and analyze datasets produced by different methods. As shown in Fig. 10, it is evident that the dataset generated by RDED [31] is overly concentrated around class centers because it originates from a pre-trained classifier. Additionally, since the pre-trained classifier cannot guarantee accurate classification results, the synthetic dataset from RDED suffers from class mixing issues. Similarly, the Minimax [7], which uses the DiT pre-trained on ImageNet-1K to generate images, also results in datasets that are overly concentrated in class centers. This concentration can be detrimental to model learning when IPC increases.
On the other hand, D4M [30] generates images using SD v1.5, which is entirely unrelated to ImageNet-1K, resulting in datasets with excessive noise and outlier samples. Although this approach can broadly cover the original input distribution, the overly vague classification boundaries hinder its effectiveness in providing useful guidance for dataset distillation tasks. In contrast to all the aforementioned methods, our approach achieves a good balance between diversity and representativeness by using SD v1.5 for localization rather than generation. Furthermore, our method can still leverage diffusion models to generate images, allowing for a mixed dataset of original and generated images, which may lead to improved performance.
8 Visualizations
To facilitate a more intuitive understanding of our methodology, we present the complete synthetic datasets for ImageWoof and ImageNette at IPC=10 as illustrated in Figs. 11 and 12, as well as the partial synthetic dataset at IPC=50, depicted in Figs. 13 and 14. Furthermore, we have visualized all images belonging to the “Golden Retriever” class at IPC=100, as shown in Fig. 15.
Figure 10: Comparison of t-SNE visualizations of features extracted by a pre-trained ResNet-18 on ImageWoof under IPC=100.
Figure 11: Visualization of ImageWoof under IPC=10.
Figure 12: Visualization of ImageNette under IPC=10.
Figure 13: Partial visualization of ImageWoof under IPC=50.
Figure 14: Partial visualization of ImageNette under IPC=50.
Figure 15: Visualization of class “Golden Retriever” under IPC=100.